A fisherman in a canoe paddles 14 miles down a stream in the same amount of time that it takes him to paddle 6 miles upstream. If the speed of the current is 2 mph, what is the speed at which the fisherman can paddle in still water?
Answer by stanbon(48558)

A fisherman in a canoe paddles 14 miles down a stream in the same amount of time that it takes him to paddle 6 miles upstream. If the speed of the current is 2 mph, what is the speed at which the fisherman can paddle in still water?
------------------------
Downstream DATA:
distance = 14 miles ; rate = b+2 mph ; time = d/r = 14/(b+2) hrs.
---------------
Upstream DATA:
distance = 6 miles ; rate = b-2 mph ; time = d/r = 6/(b-2) hrs.
---------------
EQUATION:
time up = time down
6/(b-2) = 14/(b+2)
6(b+2) = 14(b-2)
6b + 12 = 14b -28
8b = 40
b = 5 mph (speed of the boat in still water)
=============================================
Cheers,
Stan H.